OpenStudy (freemap):

1. Consider the system of equations. x-5y=2 3x-10y=11 (a) Jamar rewrote the first equation as –2x + 10y = –4. What justification did he have for doing this? (b) Jamar combined the equation in Part (a) with the second equation in the system to get x = 7. Explain Jamar’s reasoning. (c) What is the solution of the system? (d) How do you know, without graphing, that the graph of 4x – 15y = 13 passes through the point whose coordinates are the solution of the system?

OpenStudy (anonymous):

hello

OpenStudy (mathstudent55):

@freemap Compare the equation in part a with the first original equation. Original first equation: x - 5y = 2 Equation from part a: –2x + 10y = –4 What do you do to x to become -2x? What do you do to -5y to become 10y? What do you do to 2 to become -4? All questions above have the same answer. What is it?

OpenStudy (freemap):

please hold

OpenStudy (freemap):

@mathstudent55 same answer so add

OpenStudy (mathstudent55):

No. Adding won't do it. There is no single quantity you can add in those three cases that will work. Try multiplication.

OpenStudy (freemap):

Yes, multiplication because 3(5y+2)

OpenStudy (mathstudent55):

Where did you get that from?

OpenStudy (mathstudent55):

What do you do multiply x by to get -2x? What do you do multiply -5y by to get 10y? What do you do multiply 2 by to get -4? All questions have the same answer.

OpenStudy (freemap):

3(5y+2)-10y=11 5y=5 y=1

OpenStudy (mathstudent55):

You're using the substitution method, but this problem is not about substitution.

OpenStudy (freemap):

@mathstudent55 multiplying by 2

OpenStudy (freemap):

@jdoe0001

OpenStudy (freemap):

@KamiBug

OpenStudy (freemap):

@retirEEd

OpenStudy (freemap):

@rebeccaxhawaii

OpenStudy (freemap):

@jhonyy9

OpenStudy (freemap):

@jigglypuff314

OpenStudy (freemap):

I need part D @jdoe0001

OpenStudy (freemap):

I need part D @jhonyy9

jhonyy9 (jhonyy9):

ok. what is your opinon ,any idea about part D. ?

jhonyy9 (jhonyy9):

what mean eaxctly solution of a system ?

OpenStudy (jdoe0001):

right.. what is the solution you've got from part 7c?

OpenStudy (jdoe0001):

so, you're doing part D, so... we assume part 7c is already done

OpenStudy (freemap):

yes only need part D

OpenStudy (freemap):

@jhonyy9